OpenStudy (stamp):

Use the following function \[g(x)=x^3-(-3x^2)-9x-7\] On what intervals is the function decreasing?

OpenStudy (stamp):

To know what intervals the function is decreasing, we need to know g'(x).\[g'(x)=3x^2+6x-9\]

OpenStudy (stamp):

Solve the zeros for g'(x) so you can see the signs of the function by and around those points.\[3x^2+6x-9=0\]\[3(x^2+2x-3)=0\]\[(x+3)(x-1)\] Zeros are x = -3, 1

OpenStudy (stamp):

g'(-4) is +, while g'(-2) is - g'(0) is -, while g(2) is +

OpenStudy (stamp):

The function is decreasing from (-3,1)

OpenStudy (stamp):

Graph of the function confirms

OpenStudy (loser66):

Wow, impressive student.

OpenStudy (stamp):

Word

OpenStudy (loser66):

means?

OpenStudy (stamp):

word is "cool, alright"